OpenStudy (anonymous):

1. Write a proportion that can be used to find the cost of 10 note books if 3 notebooks cost$1.98 3/10 =n/$1.98 10/$1.99 =n/3 10/3 = $1.98/n 3/$1.98= 10/9 Please help i'll give you a medal I'll also fan you

OpenStudy (anonymous):

The proportion is a fraction. You can set it up as: \[\frac{price}{notebooks}\] or \[\frac{notebooks}{price}\] doesn't matter. And this fraction will be maintained. So if you know that 3 notebooks cost 1.98 that's the correct proportion, you can see it as. \[\frac{notebooks}{price}=\frac{3}{1.98}=\frac{10}{n}\]

OpenStudy (anonymous):

thank you so much.

OpenStudy (anonymous):

I just gave a medal